OpenStudy (anonymous):

find the general indefinite integral of (sqrtx^3+cubedrtx^2)

OpenStudy (anonymous):

so the integral of x^3/2 +x^2/3? cause that is 2/5 x^5/2 +3/5 x^5/3 +C

OpenStudy (anonymous):

\[\int\limits_{z}^{z}\sqrt{x}^3+\sqrt[3]{x}^2\] minus the z and x is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think it's 2x^(5/2) + 3x^(5/3) ---------------------- 5

OpenStudy (anonymous):

oh plus a constant

OpenStudy (anonymous):

same thing as i had...you just made it one fraction instead of two

OpenStudy (anonymous):

oh yeah I didn't read what you put lol took me a while cause I just learned how to do it this week